# Prediction-Powered Adaptive Shrinkage Estimation

### v2

@@ Appendix B The Correlation-Aware Unbiased Risk Estimate
Stein’s Unbiased Risk Estimate (SURE) was proposed in Charles Stein’s seminal work (1981) to study the quadratic risk in Gaussian sequence models. As a simple special case of SURE, let $Z\sim\mathcal{N}(\theta,\,\sigma^{2})$ and let $h:\mathbb{R}\to\mathbb{R}$ be an absolutely continuous function and $\mathbb{E}_{\theta}\!\left[|h^{\prime}(Z)|\right]<\infty$, then SURE is defined as
@@
with the property that $\mathbb{E}_{\theta}\!\left[\text{SURE}(h)\right]=R(h(Z),\theta)=\mathbb{E}_{\theta}[(h(Z)-\theta)^{2}]$. A proof of this argument relies on Stein’s lemma, an identity specific to Gaussian random variables (Stein, 1981). Now consider the specific linear shrinkage estimator $h_{c}(Z):=cZ+(1-c)Y$, with $c\in[0,1]$ and $Y\in\mathbb{R}$ being fixed (that is, $Y$ is a constant, or $Y$ is independent of $Z$ and we condition on $Y$). Then $\mathrm{SURE}$ takes the following form:
@@
=$\stackrel{{\scriptstyle(\star)}}{{=}}\text{CURE}(h_{c}(Z)),$
where in ()($\star$)( ⋆ ) we used the fact that in this case (with $Y$ fixed or $Y$ independent of $Z$), it holds that $\gamma=0$ so that the definition of $\mathrm{CURE}$ in (22) simplifies. This explains how CURE defined in B.1 is connected to SURE.
We make one last remark: The derivation of SURE itself requires Gaussianity. However, for linear shrinkage rules as $h_{c}(Z)$, SURE only depends on the first two moments of the distribution of $Z$ and thus is an unbiased estimator of quadratic risk under substantial generality as long as $\mathbb{E}_{\theta}[Z]=\theta$ and $\mathrm{Var}_{\theta}[Z]=\sigma^{2}$. This remark has been made by previous authors, e.g., Kou & Yang (2017); Ignatiadis & Wager (2019) and is important for the assumption-lean validity of PAS.
@@ Appendix C Details on PAS with Unknown Second Moments
[…]
@@ C.2 UniPT: Power-tuning Across Problems with Estimated Second Moments
    $\mathbb{E}_{\mathbb{P}_{\eta}}[\gamma_{j}^{2}]<\infty$, $\>\mathbb{E}_{\mathbb{P}_{\eta}}[(\tau_{j}^{2})^{2}]<\infty,\>\mathbb{E}_{\mathbb{P}_{\eta}}[(\sigma_{j}^{2})^{2}]<\infty$. [^18] Note that this condition is implied by the finite fourth-moment assumption in Proposition 5.1.
@@
[…]
@@ C.3 UniPAS: CURE and Adaptive Shrinkage with Estimated Second Moments
12:  $\triangleright$ Step 4: Adaptive shrinkage
13:  $\bar{\sigma}^{2}=\frac{1}{m}\sum_{j=1}^{m}\hat{\sigma}_{j}^{2},\>\bar{\tau}^{2}=\frac{1}{m}\sum_{j=1}^{m}\hat{\tau}_{j}^{2},\>\bar{\gamma}=\frac{1}{m}\sum_{j=1}^{m}\hat{\gamma}_{j}$
@@
18:  $\triangleright$ Now for each choice of $\omega\geq 0$, the shrinkage coefficient is determined by $\frac{\omega}{\omega+\check{\sigma}^{2}}$
19:  $\triangleright$ The CURE calculation in Equation 26, which still depends on the sample-based estimators $\dot{\sigma}_{j}^{2},\dot{\gamma}_{j}$
20:  $\hat{\omega}=\texttt{{get\_shrink\_param}}((\hat{\theta}_{j}^{\mathrm{UPT}})_{j=1}^{m},(\tilde{Z}^{f}_{j})_{j=1}^{m},(\check{\sigma}_{j}^{2})_{j=1}^{m},(\dot{\sigma}_{j}^{2})_{j=1}^{m},(\dot{\gamma}_{j})_{j=1}^{m})$
@@
$\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\hat{\boldsymbol{\theta}}^{\mathrm{UPAS}}\Big{)}\leq\inf_{\omega\geq 0}\left\{\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\boldsymbol{\hat{\theta}}^{\mathrm{UPAS}}_{\omega}\Big{)}\right\}+o(1)\quad\text{ as }\;m\to\infty,$
@@
$\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\hat{\boldsymbol{\theta}}^{\mathrm{UPAS}}\Big{)}\leq\min\left\{\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\boldsymbol{\tilde{Z}^{f}}\Big{)},\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\hat{\boldsymbol{\theta}}^{\mathrm{UPT}}\Big{)}\right\}+o(1)\quad\text{ as }\;m\to\infty.$
@@ Appendix D Other Baseline Shrinkage Estimators
where $\text{where }\quad\omega_{j}$ $:=\omega/(\omega+\tilde{\sigma}^{2}_{j}),\quad\tilde{\sigma}^{2}_{j}:=\text{Var}_{\eta_{j}}[\bar{Y}_{j}]=\sigma_{j}^{2}/n_{j}.$
@@
where $\text{where }\quad\hat{\omega}\in\operatorname*{arg\,min}_{\omega\geq 0}\frac{1}{m}\sum_{j=1}^{m}\mathrm{CURE}(\hat{\theta}_{j,\omega}^{\text{Shrink}}).$ (28)
@@
5:  $\hat{\omega}=\texttt{{get\_shrink\_param}}((\bar{Y}_{j})_{j=1}^{m},(\tilde{Z}^{f}_{j})_{j=1}^{m},(\tilde{\sigma}_{j}^{2})_{j=1}^{m})$triangleright use Eq. (D.1)
@@
7:    $\tilde{\sigma}_{j}^{2}=\texttt{{get\_pt\_var}}(\hat{\theta}_{j}^{\mathrm{PT}})$triangleright (Eq. (14))
@@
where $\text{where }\quad\omega_{j}$ $:=\omega/(\omega+\tilde{\sigma}^{2}_{j}),\quad\tilde{\sigma}^{2}_{j}:=\mathrm{Var}_{\eta_{j}}[\hat{\theta}^{\text{PT}}_{j}].$
@@
Xie et al. (2012) proposed the following unbiased risk estimate to optimize $\omega$ for this estimator. Note that even though the group mean is also correlated with each PT estimator, we still denote the following SURE instead of CURE following the nomenclature in Xie et al. (2012).
@@ E.2 Amazon Review Ratings Dataset
The Amazon Fine Food Reviews dataset, provided by the Stanford Network Analysis Project (SNAP; SNAP (2014)) on Kaggle, [^21] https://www.kaggle.com/datasets/snap/amazon-fine-food-reviews comes in a clean format. We group reviews by their ProductID. For each review, we concatenate the title and body text to form the covariate, while the response is the reviewer’s score/rating (1 to 5 stars). Here’s a sample review:
@@
The Bidirectional Encoder Representations from Transformers (BERT) model is a widely adopted language model for many NLP tasks including text classification (Devlin et al., 2019). However, pre-training BERT from scratch is time-consuming and requires large amounts of data. We therefore use the bert-base-multilingual-uncased-sentiment model [^22] https://huggingface.co/nlptown/bert-base-multilingual-uncased-sentiment from Town (2023) as the base model, denoted as BERT-base. BERT-base is pre-trained on general product reviews (not exclusive to Amazon) in six languages. It achieves 67.5% prediction accuracy on a validation set of 100 products ($\sim$46k reviews).
Then, we further fine-tune it on the held-out review data, that is, reviews outside the top 200 products, for 2 full epochs. The fine-tuning is done using Hugging Face’s transformers library (Wolf, 2019). After fine-tuning, the BERT-tuned model achieves 78.8% accuracy on the same validation set.
@@ E.3 Spiral Galaxy Fractions (Galaxy Zoo 2)
The Galaxy Zoo 2 (GZ2) project [^23] https://data.galaxyzoo.org/ contains a large collection of human-annotated classification results for galaxy images from SDSS. However, instead of having a single dataframe, GZ2 has many different tables—each for subsets of the SDSS raw data. We begin with a particular subset of 239,696 images with metadata drawn from Hart et al. (2016). Our data cleaning pipeline is inspired by Lin et al. (2021), which removes missing data and relabels the class name of each galaxy image to a more readable format:
@@
After partitioning the images based on WVT_BIN, [^25] In the Galaxy Zoo 2 dataset, WVT_BIN denotes Voronoi bins constructed based on each galaxy’s intrinsic size and absolute magnitude. The motivation and implementation of this binning strategy are detailed in Hart et al. (2016), who justify such partitioning—aligned with our compound mean estimation setup—by noting that spiral arm morphology exhibits systematic dependencies on stellar mass and related intrinsic properties. we consider only the top $m=100$ partitions based on the cutoff that each problem should have $\geq 150$ images (many partitions have very few galaxy images in them), for the same reason as in the Amazon Review dataset. Finally, we have a total of $\sim$100k images as covariates (either $X_{ij}$ or $\tilde{X}_{ij}$) for our problem.
@@
We employ the ResNet50 architecture (He et al., 2016), utilizing the pre-trained model from torchvision initially trained on ImageNet (Deng et al., 2009). To tailor the model to our task, we fine-tune it on $\sim$50k images excluded from the top $m$ problems. The model is trained to classify galaxies into eight categories, later condensed into a binary spiral/non-spiral classification for prediction. We use a batch size of 256 and Adam optimizer (Kingma & Ba, 2015) with a learning rate of 1e-3. After 20 epochs, the model achieves 87% training accuracy and 83% test accuracy. Despite these promising results, Table 3 indicates that the predictions still require debiasing for accurate estimation.
@@ E.4 Benchmarking in real-world datasets
    Following evaluation methodology in existing PPI literature, e.g., (Angelopoulos et al., 2023), we first calculate the mean of all responses for each problem and treat it as the pseudo ground-truth, i.e., $\dot{\theta}_{j}:=\frac{1}{T_{j}}\sum_{i}\dot{Y}_{ij}$.
@@
    $\widehat{\text{MSE}}_{K}(\boldsymbol{\hat{\theta}})$ $:=\frac{1}{K}\sum_{k=1}^{K}\left(\frac{1}{m}\sum_{j=1}^{m}(\hat{\theta}_{j}^{k}-\dot{\theta}_{j})^{2}\right),\quad\text{SE}_{K}(\boldsymbol{\hat{\theta}}):=\frac{1}{\sqrt{K}}\sqrt{\frac{1}{K-1}\sum_{k=1}^{K}\left(\frac{1}{m}\sum_{j=1}^{m}(\hat{\theta}_{j}^{k}-\dot{\theta}_{j})^{2}-\widehat{\text{MSE}}_{K}(\boldsymbol{\hat{\theta}})\right)}.$
@@ Appendix F Proofs of Theoretical Results
The proof in this step is directly adapted from Theorem 5.1 in Xie et al. (2012) and generalizes to non-Gaussian data. With some algebraic manipulation, we can further decompose
@@
For the second term ()(**)( ∗ ∗ ), we start by arguing conditionally on $\boldsymbol{\eta}$, which implies in particular that we may treat all the $\tilde{\sigma}^{2}_{j}$ as fixed. It is thus without loss of generality to assume that $\tilde{\sigma}^{2}_{1}\leq...\leq\tilde{\sigma}^{2}_{m}$ (by first sorting problems according to the value of $\tilde{\sigma}^{2}_{j}$). Then, since $\omega_{j}$ is monotonic function of $\tilde{\sigma}_{j}^{2}$ for any fixed $\omega\geq 0$, we have $1\geq\omega_{1}\geq...\geq\omega_{m}\geq 0$. The following inequality follows:
@@
Let $M_{k}:=\sum_{j=1}^{k}(\big{(}\hat{\theta}^{\text{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\text{PT}}_{j}\theta_{j}-\tilde{\sigma}^{2}_{j})$, it is easy to see that $\{M_{k}\}_{k=1}^{m}$ forms a martingale conditional on $\boldsymbol{\eta}$. Therefore, by a standard $L^{2}$ maximal inequality (ref. Theorem 4.4.6 in Durrett (2019)), we have
@@
Noting that the right hand side converges to 00 as $m\to\infty$, and recalling the definition of $\omega_{*}$, we prove the desired result
@@
So $\lambda^{*}_{m}=N_{m}^{*}/D_{m}^{*}$ and $\hat{\lambda}=\hat{N}_{m}/\hat{D}_{m}$. We denote $\Delta N_{m}=\hat{N}_{m}-N_{,}^{*}=m^{-1}\sum_{j=1}^{m}U_{j}$ where $U_{j}=n_{j}^{-1}(\hat{\gamma}_{j}-\gamma_{j})$. Note that $\mathbb{E}_{\mathbb{P}_{\eta}}[U_{j}]=0$, and by assumption 2 and 3 we know that $\mathrm{Var}_{\mathbb{P}_{\eta}}[U_{j}]=\mathbb{E}_{\mathbb{P}_{\eta}}[n_{j}^{-2}\mathrm{Var}_{\eta_{j}}[\hat{\gamma}_{j}]]$ is bounded (say $<V_{U}<\infty$), we then have
@@
[…]
@@
[…]
@@
[…]
@@
$=(2A_{m}\xi_{m}-2B_{m})(\hat{\lambda}_{\mathrm{clip}}-\lambda^{*}_{\mathrm{clip},m})$
[…]
@@
We provide a high-level sketch of the proof in Figure 7. Our main goal is to establish that $\widehat{\text{CURE}}-\ell_{m}\xrightarrow{L^{1}}0$ uniformly in $\omega$ as $m\to\infty$. To achieve this, we introduce two intermediate estimators—$\text{CURE}^{\prime}$ and $\text{CURE}^{\prime\prime}$—that serve as bridges between $\widehat{\text{CURE}}$ and $\ell_{m}$. The proof proceeds by showing that each consecutive pair of estimators is asymptotically close, which allows us to conclude the overall result via repeated applications of the triangle inequality.
$\widehat{\text{CURE}}$q. (26)$\text{CURE}^{\prime}$q. (40)$\text{CURE}^{\prime\prime}$q. (41)$\ell_{m}$q. (9)Lemma F.7Lemma F.5Theorem F.4
@@
By construction, $\text{CURE}^{\prime}$ and $\text{CURE}^{\prime\prime}$ only differ in their use of $\dot{\sigma}_{j}^{2}$ versus $\ddot{\sigma}_{j}^{2}$ (similarly for $\dot{\gamma}_{j}$ v.s.$\ddot{\gamma}_{j}$). We can thus decompose their difference as
@@
These uniform $L^{2}$ convergence conditions are sufficient to make sure the expectation of Section F.7 converges to 00, and we thus prove our lemma. ∎
@@
[…]

### v3

@@ Appendix B The Correlation-Aware Unbiased Risk Estimate
Stein’s Unbiased Risk Estimate (SURE) was proposed in Charles Stein’s seminal work (stein1981estimation) to study the quadratic risk in Gaussian sequence models. As a simple special case of SURE, let $Z\sim\mathcal{N}(\theta,\,\sigma^{2})$ and let $h:\mathbb{R}\to\mathbb{R}$ be an absolutely continuous function and $\mathbb{E}_{\theta}\!\left[|h^{\prime}(Z)|\right]<\infty$, then SURE is defined as
@@
with the property that $\mathbb{E}_{\theta}\!\left[\text{SURE}(h)\right]=R(h(Z),\theta)=\mathbb{E}_{\theta}[(h(Z)-\theta)^{2}]$. A proof of this argument relies on Stein’s lemma, an identity specific to Gaussian random variables (stein1981estimation). Now consider the specific linear shrinkage estimator $h_{c}(Z):=cZ+(1-c)Y$, with $c\in[0,1]$ and $Y\in\mathbb{R}$ being fixed (that is, $Y$ is a constant, or $Y$ is independent of $Z$ and we condition on $Y$). Then $\mathrm{SURE}$ takes the following form:
@@
$\stackrel{{\scriptstyle(\star)}}{{=}}\text{CURE}(h_{c}(Z)),$
where in $(\star)$ we used the fact that in this case (with $Y$ fixed or $Y$ independent of $Z$), it holds that $\gamma=0$ so that the definition of $\mathrm{CURE}$ in (22) simplifies. This explains how CURE defined in B.1 is connected to SURE.
We make one last remark: The derivation of SURE itself requires Gaussianity. However, for linear shrinkage rules as $h_{c}(Z)$, SURE only depends on the first two moments of the distribution of $Z$ and thus is an unbiased estimator of quadratic risk under substantial generality as long as $\mathbb{E}_{\theta}[Z]=\theta$ and $\mathrm{Var}_{\theta}[Z]=\sigma^{2}$. This remark has been made by previous authors, e.g., kou2017optimal; ignatiadis2019covariatepowered and is important for the assumption-lean validity of PAS.
@@ Appendix C Details on PAS with Unknown Second Moments
Throughout this section and the proof details in Appendix F, we use $\xrightarrow{P}$ for convergence in probability and $\xrightarrow{L^{p}}$ for $L^{p}$ convergence of random variables. We use the “Little-o” notation $o(1)$ for any term that vanishes to zero as $m\to\infty$. Similarly, for a sequence of random variables $X_{m}$, we use $X_{m}=o_{P}(1)$ if $X_{m}$ converges to zero in probability, and $X_{m}=O_{P}(1)$ if $X_{m}$ is bounded in probability. All stochastic order relations are understood to hold as $m\to\infty$.
@@ C.2 UniPT: Power-tuning Across Problems with Estimated Second Moments
    $\mathbb{E}_{\mathbb{P}_{\eta}}[\gamma_{j}^{2}]<\infty$, $\mathbb{E}_{\mathbb{P}_{\eta}}[(\tau_{j}^{2})^{2}]<\infty,\>\mathbb{E}_{\mathbb{P}_{\eta}}[(\sigma_{j}^{2})^{2}]<\infty$. [^18] Note that this condition is implied by the finite fourth-moment assumption in Proposition 5.1.
@@
$\frac{1}{m}\sum_{j=1}^{m}\mathrm{Var}_{\eta_{j}}[\hat{\theta}_{j}^{\mathrm{UPT}}]-\min_{\lambda^{\prime}\in[0,1]}\frac{1}{m}\sum_{j=1}^{m}\mathrm{Var}_{\eta_{j}}[\hat{\theta}_{j,\lambda^{\prime}}]\xrightarrow{P}\,0\quad\text{as }\,m\to\infty,$
@@ C.3 UniPAS: CURE and Adaptive Shrinkage with Estimated Second Moments
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Appendix D Other Baseline Shrinkage Estimators
$\text{where }\quad\omega_{j}$ $:=\omega/(\omega+\tilde{\sigma}^{2}_{j}),\quad\tilde{\sigma}^{2}_{j}:=\text{Var}_{\eta_{j}}[\bar{Y}_{j}]=\sigma_{j}^{2}/n_{j}.$
@@
$\text{where }\quad\hat{\omega}\in\operatorname*{arg\,min}_{\omega\geq 0}\frac{1}{m}\sum_{j=1}^{m}\mathrm{CURE}(\hat{\theta}_{j,\omega}^{\text{Shrink}}).$ (28)
@@
5:  $\hat{\omega}=\texttt{{get\_shrink\_param}}((\bar{Y}_{j})_{j=1}^{m},(\tilde{Z}^{f}_{j})_{j=1}^{m},(\tilde{\sigma}_{j}^{2})_{j=1}^{m})$ $\triangleright$ use Eq. (D.1)
@@
7:  $\tilde{\sigma}_{j}^{2}=\texttt{{get\_pt\_var}}(\hat{\theta}_{j}^{\mathrm{PT}})$ $\triangleright$ (Eq. (14))
@@
$\text{where }\quad\omega_{j}$ $:=\omega/(\omega+\tilde{\sigma}^{2}_{j}),\quad\tilde{\sigma}^{2}_{j}:=\mathrm{Var}_{\eta_{j}}[\hat{\theta}^{\text{PT}}_{j}].$
@@
xie2012sure proposed the following unbiased risk estimate to optimize $\omega$ for this estimator. Note that even though the group mean is also correlated with each PT estimator, we still denote the following SURE instead of CURE following the nomenclature in xie2012sure.
@@ E.2 Amazon Review Ratings Dataset
The Amazon Fine Food Reviews dataset, provided by the Stanford Network Analysis Project (SNAP; amazon_fine_food_reviews) on Kaggle, [^21] https://www.kaggle.com/datasets/snap/amazon-fine-food-reviews comes in a clean format. We group reviews by their ProductID. For each review, we concatenate the title and body text to form the covariate, while the response is the reviewer’s score/rating (1 to 5 stars). Here’s a sample review:
@@
The Bidirectional Encoder Representations from Transformers (BERT) model is a widely adopted language model for many NLP tasks including text classification (devlin2019bert). However, pre-training BERT from scratch is time-consuming and requires large amounts of data. We therefore use the bert-base-multilingual-uncased-sentiment model [^22] https://huggingface.co/nlptown/bert-base-multilingual-uncased-sentiment from nlp_town_2023 as the base model, denoted as BERT-base. BERT-base is pre-trained on general product reviews (not exclusive to Amazon) in six languages. It achieves 67.5% prediction accuracy on a validation set of 100 products ($\sim$46k reviews).
Then, we further fine-tune it on the held-out review data, that is, reviews outside the top 200 products, for 2 full epochs. The fine-tuning is done using Hugging Face’s transformers library (wolf2019huggingface). After fine-tuning, the BERT-tuned model achieves 78.8% accuracy on the same validation set.
@@ E.3 Spiral Galaxy Fractions (Galaxy Zoo 2)
The Galaxy Zoo 2 (GZ2) project [^23] https://data.galaxyzoo.org/ contains a large collection of human-annotated classification results for galaxy images from SDSS. However, instead of having a single dataframe, GZ2 has many different tables—each for subsets of the SDSS raw data. We begin with a particular subset of 239,696 images with metadata drawn from hart2016galaxy. Our data cleaning pipeline is inspired by lin2021galaxy, which removes missing data and relabels the class name of each galaxy image to a more readable format:
@@
After partitioning the images based on WVT_BIN, [^25] In the Galaxy Zoo 2 dataset, WVT_BIN denotes Voronoi bins constructed based on each galaxy’s intrinsic size and absolute magnitude. The motivation and implementation of this binning strategy are detailed in hart2016galaxy, who justify such partitioning—aligned with our compound mean estimation setup—by noting that spiral arm morphology exhibits systematic dependencies on stellar mass and related intrinsic properties. we consider only the top $m=100$ partitions based on the cutoff that each problem should have $\geq 150$ images (many partitions have very few galaxy images in them), for the same reason as in the Amazon Review dataset. Finally, we have a total of $\sim$100k images as covariates (either $X_{ij}$ or $\tilde{X}_{ij}$) for our problem.
@@
We employ the ResNet50 architecture (he2016deep), utilizing the pre-trained model from torchvision initially trained on ImageNet (deng2009imagenet). To tailor the model to our task, we fine-tune it on $\sim$50k images excluded from the top $m$ problems. The model is trained to classify galaxies into eight categories, later condensed into a binary spiral/non-spiral classification for prediction. We use a batch size of 256 and Adam optimizer (kingma2014adam) with a learning rate of 1e-3. After 20 epochs, the model achieves 87% training accuracy and 83% test accuracy. Despite these promising results, Table 3 indicates that the predictions still require debiasing for accurate estimation.
@@ E.4 Benchmarking in real-world datasets
    Following evaluation methodology in existing PPI literature, e.g., (angelopoulos2023prediction), we first calculate the mean of all responses for each problem and treat it as the pseudo ground-truth, i.e., $\dot{\theta}_{j}:=\frac{1}{T_{j}}\sum_{i}\dot{Y}_{ij}$.
@@
[…]
@@ Appendix F Proofs of Theoretical Results
The proof in this step is directly adapted from Theorem 5.1 in xie2012sure and generalizes to non-Gaussian data. With some algebraic manipulation, we can further decompose
@@
For the second term ()(**), we start by arguing conditionally on $\boldsymbol{\eta}$, which implies in particular that we may treat all the $\tilde{\sigma}^{2}_{j}$ as fixed. It is thus without loss of generality to assume that $\tilde{\sigma}^{2}_{1}\leq...\leq\tilde{\sigma}^{2}_{m}$ (by first sorting problems according to the value of $\tilde{\sigma}^{2}_{j}$). Then, since $\omega_{j}$ is monotonic function of $\tilde{\sigma}_{j}^{2}$ for any fixed $\omega\geq 0$, we have $1\geq\omega_{1}\geq...\geq\omega_{m}\geq 0$. The following inequality follows:
@@
Let $M_{k}:=\sum_{j=1}^{k}(\big(\hat{\theta}^{\text{PT}}_{j}\big)^{2}-\hat{\theta}^{\text{PT}}_{j}\theta_{j}-\tilde{\sigma}^{2}_{j})$, it is easy to see that $\{M_{k}\}_{k=1}^{m}$ forms a martingale conditional on $\boldsymbol{\eta}$. Therefore, by a standard $L^{2}$ maximal inequality (ref. Theorem 4.4.6 in durrett2019probability), we have
@@
Noting that the right hand side converges to 0 as $m\to\infty$, and recalling the definition of $\omega_{*}$, we prove the desired result
@@
So $\lambda^{*}_{m}=N_{m}^{*}/D_{m}^{*}$ and $\hat{\lambda}=\hat{N}_{m}/\hat{D}_{m}$. We denote $\Delta N_{m}=\hat{N}_{m}-N_{,}^{*}=m^{-1}\sum_{j=1}^{m}U_{j}$ where $U_{j}=n_{j}^{-1}(\hat{\gamma}_{j}-\gamma_{j})$. Note that $\mathbb{E}_{\mathbb{P}_{\eta}}[U_{j}]=0$, and by assumption 2 and 3 we know that $\mathrm{Var}_{\mathbb{P}_{\eta}}[U_{j}]=\mathbb{E}_{\mathbb{P}_{\eta}}[n_{j}^{-2}\mathrm{Var}_{\eta_{j}}[\hat{\gamma}_{j}]]$ is bounded (say $V_{U}<\infty$), we then have
@@
So $\Delta N_{m}\xrightarrow{L^{2}}0$, which implies $\Delta N_{m}\xrightarrow{P}0$. Thus, $\hat{N}_{m}-N_{m}^{*}\xrightarrow{P}0$. An identical proof will also give us $\hat{D}_{m}-D_{m}^{*}\xrightarrow{P}0$ for the denominator terms.
@@
$\hat{\lambda}_{\mathrm{clip}}-\lambda^{*}_{\mathrm{clip},m}=\mathrm{\mathrm{clip}}(\hat{\lambda},[0,1])-\mathrm{\mathrm{clip}}(\lambda^{*}_{m},[0,1])\xrightarrow{P}0.$
@@
We want to show that $V_{m}(\hat{\lambda}_{\mathrm{clip}})-V_{m}(\lambda^{*}_{\mathrm{clip},m})\xrightarrow{P}0$. The derivative of $V_{m}(\lambda)$ is $V^{\prime}_{m}(\lambda)=2A_{m}\lambda-2B_{m}$. By the Mean Value Theorem, for some $\xi_{m}$ between $\hat{\lambda}_{\mathrm{clip}}$ and $\lambda^{*}_{\mathrm{clip},m}$:
@@
$(2A_{m}\xi_{m}-2B_{m})(\hat{\lambda}_{\mathrm{clip}}-\lambda^{*}_{\mathrm{clip},m})$
Since $\hat{\lambda}_{\mathrm{clip}}$ and $\lambda^{*}_{\mathrm{clip},m}$ are both in $[0,1]$, $\xi_{m}$ is also in $[0,1]$. The terms $A_{m}=D_{m}^{*}$ and $B_{m}=N_{m}^{*}$ are averages of $m$ independent terms with uniformly bounded variances (under Assumption 4 of Lemma C.2). Thus, by Chebyshev’s inequality, $A_{m}=O_{P}(1)$ and $B_{m}=O_{P}(1)$ (i.e., they are bounded in probability). Since $\xi_{m}\in[0,1]$, the term $(2A_{m}\xi_{m}-2B_{m})$ is also $O_{P}(1)$. Let $\Delta\lambda_{m}=\hat{\lambda}_{\mathrm{clip}}-\lambda^{*}_{\mathrm{clip},m}$. From Lemma C.2, we have $\Delta\lambda_{m}\xrightarrow{P}0$. Therefore,
$V_{m}(\hat{\lambda}_{\mathrm{clip}})-V_{m}(\lambda^{*}_{\mathrm{clip},m})=(2A_{m}\xi_{m}-2B_{m})\Delta\lambda_{m}\xrightarrow{P}0.$
@@
[…]
@@
[…]
@@
These uniform $L^{2}$ convergence conditions are sufficient to make sure the expectation of Section F.7 converges to 0, and we thus prove our lemma. ∎
@@
$\bar{\sigma}^{2}\xrightarrow{P}\mathbb{E}_{\mathbb{P_{\eta}}}[\sigma^{2}_{j}]=:\mu_{\sigma^{2}},\quad\bar{\tau}^{2}\xrightarrow{P}\mathbb{E}_{\mathbb{P_{\eta}}}[\tau^{2}_{j}]=:\mu_{\tau^{2}},\quad\bar{\gamma}\xrightarrow{P}\mathbb{E}_{\mathbb{P_{\eta}}}[\gamma_{j}]=:\mu_{\gamma},$